### v2

# The heat trace for domains with curved corners

Abstract.
We derive the short-time expansion of the heat trace on planar curvilinear polygons through order $t^{1/2}$, with both Dirichlet and Neumann boundary conditions. For such domains we show that the coefficient of $t^{1/2}$ in the expansion splits into a boundary integral of $\kappa^{2}$ and a sum of local corner contributions, one for each vertex. Each curved corner contribution depends only on the interior angle $\alpha$ and on the limiting curvatures $\kappa_{\pm}$ on the adjacent sides. For Dirichlet boundary conditions, using a conformal model and a parametrix construction on the sector heat space, we express this contribution in the form $c_{1/2}(\alpha)\,r_{0}(\alpha,\kappa_{+},\kappa_{-})$, where $c_{1/2}(\alpha)$ is given by the Hadamard finite part of an explicit trace over the exact sector. For right angles we compute $c_{1/2}(\pi/2)=1/(16\sqrt{\pi})$, which yields a closed formula for the $\sqrt{t}$ coefficient for domains with right-angled corners. Combined with previous work on the coefficients of $t^{-1}$, $t^{-1/2}$ and $t^{0}$ for such domains, this determines the remaining local heat trace coefficient for planar curvilinear polygons up to order $t^{1/2}$. As an inverse spectral application we extend a result in the literature by showing that any admissible curvilinear polygon that is Dirichlet isospectral to a polygon must itself be a polygon with straight sides.
1. Introduction
A classical problem in spectral geometry is to determine what geometric information is encoded in the spectrum of the Laplacian on a Riemannian manifold [8]. This is frequently approached using heat trace methods. Specifically, one examines the short-time asymptotic expansion of the trace of the heat kernel. This trace is spectrally determined, and therefore the coefficients in the asymptotic expansion are spectral invariants – geometric quantities determined by the spectrum. In 1967, McKean and Singer computed the first three heat invariants for a smoothly bounded domain in $\mathbb{R}^{2}$ [7]. Using these, they showed that the area, perimeter, and Euler characteristic of such a domain are all spectral invariants. It is natural to ask whether one can deduce similar information for a less smooth domain, and indeed this has been studied in depth [2, 11, 10].
We now explain the setting of this paper. Let $\Omega\subseteq\mathbb{R}^{2}$ be a curvilinear polygon, with piecewise smooth boundary, smooth except for a finite number $n$ of corners $P_{j}$ of interior angles $\alpha_{j}>0$, $j=1,\ldots,n$. Let $\kappa$ be the inward-pointing curvature of the boundary, and let $\gamma(s)$ be an arc-length parametrization of the boundary, oriented counterclockwise. For each $j$, let $\kappa_{j,\pm}$ be the limits of $\kappa(\gamma(s))$ as $s$ approaches $\gamma^{-1}(P_{j})$ from above and below respectively.
It is well known that the trace of the heat kernel on $\Omega$ with either Dirichlet or Neumann boundary conditions, $H^{D}_{\Omega}(t)$ or $H^{N}_{\Omega}(t)$ respectively, has an asymptotic expansion as $t\to 0$, beginning with a $t^{-1}$ term. The $t^{-1}$, $t^{-1/2}$, and $t^{0}$ terms of this expansion have been computed, first in the exact polygonal setting in unpublished work of Ray, as cited in [2, 12], and then extended to the general curvilinear setting in [10]. The next local heat invariant is the coefficient of $t^{1/2}$; it is the first term beyond the regime treated in [2, 11, 10] and the first coefficient that couples boundary curvature to the corner geometry. In [4], using results such as [1], this $t^{1/2}$ coefficient is computed in the “straight corners” case, where a neighborhood of each corner $P_{j}$ is isometric to an exact sector. Without this assumption, the curvature of the sides may interact with the angle at the corner. As we show, this interaction produces a term at order $t^{1/2}$.
Our first main result identifies this $t^{1/2}$ coefficient, showing that it splits into a boundary integral of $\kappa^{2}$ and a sum of local corner terms, one for each vertex. It thus completes the description of the heat trace asymptotics up to order $t^{1/2}$ for planar curvilinear polygons. We also obtain this result with Neumann boundary conditions, which we state in Proposition 1.
Theorem 1.
With notation as above, the Dirichlet heat trace $H^{D}_{\Omega}(t)$ has an expansion as $t\to 0$ given by
(1) $H^{D}_{\Omega}(t)=\frac{|\Omega|}{4\pi t}-\frac{|\partial\Omega|}{8\sqrt{\pi t}}+\frac{1}{12\pi}\Big(\int_{\partial\Omega}\kappa\,ds+\sum_{j=1}^{n}\frac{\pi^{2}-\alpha_{j}^{2}}{2\alpha_{j}}\Big)\\ +\sqrt{t}\Big(\frac{1}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds+\sum_{j=1}^{n}\mathcal{C}_{1/2}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})\Big)+O(t\log t),$
where $\mathcal{C}_{1/2}$ is a function only of the angle $\alpha$ and the limit curvatures $\kappa_{\pm}$.
We can say substantially more about $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})$ whenever $\alpha\neq\pi$. Define
(2) $r_{0}=r_{0}(\alpha,\kappa_{+},\kappa_{-})=\frac{1}{2}|\csc\alpha|\sqrt{\kappa_{+}^{2}+\kappa_{-}^{2}+2\kappa_{+}\kappa_{-}\cos\alpha}.$
Theorem 2.
For $\alpha\neq\pi$, the function $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})$ has the form
(3) $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})=c_{1/2}(\alpha)r_{0}(\alpha,\kappa_{+},\kappa_{-})$
where $c_{1/2}(\alpha)$ is a function depending only on $\alpha$, whose general expression is given by a Hadamard renormalized trace on an exact sector:
(4) $c_{1/2}(\alpha)$ $=-\,\operatorname*{fp}\iint_{(0,\infty)\times(0,\alpha)}\int_{0}^{1}(1-\sigma)^{-1}\sigma^{-1/2}\iint_{(0,\infty)\times(0,\alpha)}$
$\quad a\left(\frac{R}{\sqrt{1-\sigma}},\theta,\frac{\sqrt{\sigma}\,\rho}{\sqrt{1-\sigma}},\phi\right)\,b\left(\rho,\phi,\frac{R}{\sqrt{\sigma}},\theta\right)\,\,\rho\,d\rho\,d\phi\,\,R\,dR\,d\theta\,d\sigma.$
Here $a$ and $b$ are functions related to the exact heat kernel on a sector of angle $\alpha$, explicitly given by (43) and (58), and $\operatorname*{fp}$ denotes the Hadamard finite part at $R=\infty$ of the improper integral in $R$.
Moreover, $c_{1/2}(\alpha)$ has the special value
(5) $c_{1/2}\left(\frac{\pi}{2}\right)=\frac{1}{16\sqrt{\pi}}.$
The factorization (3) shows that, away from $\alpha=\pi$, the corner contribution at order $t^{1/2}$ splits cleanly into an angular factor $c_{1/2}(\alpha)$ and a curvature factor $r_{0}(\alpha,\kappa_{+},\kappa_{-})$, so that all dependence on the limiting curvatures is encoded in a single scalar quantity. This theorem is not just an integral identity, but a structural description of how corner curvature enters the expansion.
Remark.
A few comments:
  • The $O(t\log t)$ error in (1) is not necessarily optimal, and in fact we expect that it should be $O(t)$.
  • Each of $a$ and $b$ is a function only of $\alpha$ and the variables of integration, see (43), (58). Thus (4) is indeed independent of $\kappa_{\pm}$.
  • For inverse spectral results, it is crucial that (5) is nonzero. We can only verify this explicitly for $\alpha=\pi/2$. It would be of significant interest to find out for which $\alpha$ we have $c_{1/2}(\alpha)\neq 0$.
  • Theorem 2 does not work for $\alpha=\pi$ because the conformal model that we use breaks down in that case, as can be seen from the fact that $r_{0}$ is undefined when $\alpha=\pi$. So the case $\alpha=\pi$, where the boundary has an abrupt change in second derivative, remains open.
  • In (1), the corner contributions to the coefficients of $t^{-1}$ and $t^{0}$ depend only on the interior angles $\alpha_{j}$; the one-sided curvatures $\kappa_{j,\pm}$ do not appear. Theorem 2 shows that the first place where the corner contribution becomes sensitive to $(\kappa_{j,+},\kappa_{j,-})$ is at order $t^{1/2}$.
1.0.1. The Neumann case
Theorem 1 may be generalized, via identical arguments, to the case of the Neumann heat trace $H^{N}_{\Omega}(t)$. The first four terms of the Neumann heat trace expansion in the straight-corners setting are given in [4]. We obtain
Proposition 1.
Under the same assumptions as in the Dirichlet case,
(6) $H^{N}_{\Omega}(t)=\frac{|\Omega|}{4\pi t}+\frac{|\partial\Omega|}{8\sqrt{\pi t}}+\frac{1}{12\pi}\Big(\int_{\partial\Omega}\kappa\,ds+\sum_{j=1}^{n}\frac{\pi^{2}-\alpha_{j}^{2}}{2\alpha_{j}}\Big)\\ +\sqrt{t}\Big(\frac{5}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds+\sum_{j=1}^{n}\mathcal{C}^{N}_{1/2}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})\Big)+O(t\log t).$
It is likely that a theorem of the form Theorem 2 holds in the Neumann case as well. However, there are some complications in the arguments in section 3 that make it more difficult to prove.
1.1. Inverse spectral applications
In [4], Enciso and Gómez-Serrano prove that any curvilinear polygon with straight corners is in fact isospectral to a polygon. They require straight corners because of a lack of a formula in the curved corner case. Using the description of the $t^{1/2}$ coefficient above, we can strengthen their result.
Definition.
A curvilinear polygon $\Omega$ is admissible if, for each corner $P_{j}$ with opening angle $\alpha_{j}$, either
$c_{1/2}(\alpha_{j})>0\quad\text{or}\quad\kappa_{j,+}=\kappa_{j,-}=0.$
Proposition 2.
Suppose that an admissible curvilinear polygon $\Omega$ is Dirichlet isospectral to a polygon. Then $\Omega$ is a polygon.
Proof.
For a polygon, the $t^{1/2}$ term in the heat trace vanishes. So it must vanish for $\Omega$. But under our assumptions, by Theorems 1 and 2, the $t^{1/2}$ coefficient in $H^{D}_{\Omega}(t)$ is given by
(7) $\frac{1}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds+\sum_{j=1}^{n}c_{1/2}(\alpha_{j})r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})$
with $r_{0}$ defined in (2). Note that $\int_{\partial\Omega}\kappa^{2}\,ds\geq 0$. Moreover $r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})\geq 0$ by definition, and (for $\alpha_{j}\neq\pi$) one has $r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})=0$ if and only if $\kappa_{j,+}=\kappa_{j,-}=0$. By admissibility, whenever $r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})>0$ we necessarily have $c_{1/2}(\alpha_{j})>0$. Hence every term in (7) is non-negative. Since the $t^{1/2}$ coefficient vanishes, each term must vanish: $\int_{\partial\Omega}\kappa^{2}\,ds=0$ and $r_{0}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})=0$ for all $j$. Therefore $\kappa\equiv 0$ on each smooth edge and $\kappa_{j,+}=\kappa_{j,-}=0$ at every corner, so $\partial\Omega$ is piecewise linear and $\Omega$ is a polygon. ∎
Remark.
The admissibility hypothesis is used only to ensure that whenever a corner has nonzero curvature on at least one side, the corresponding coefficient satisfies $c_{1/2}(\alpha_{j})>0$. The alternative condition $\kappa_{j,+}=\kappa_{j,-}=0$ requires only vanishing one-sided curvature limits; it does not assume that the boundary is straight in a full neighborhood of $P_{j}$.
Remark.
Observe from (5) that any polygon where all non-straight corners are right angles is admissible.
Together with the known coefficients of order $t^{-1}$, $t^{-1/2}$ and $t^{0}$, Theorems 1 and 2 show that, for planar curvilinear polygons, all local heat invariants up to order $t^{1/2}$ can be expressed in explicit geometric terms and used in inverse spectral arguments of this type.
Remark (A sign condition that would enlarge the admissible class).
In our argument, admissibility is used only through the implication
$(\kappa_{j,+},\kappa_{j,-})\neq(0,0)\quad\Longrightarrow\quad c_{1/2}(\alpha_{j})>0$
at each corner $P_{j}$. Any structural information on the sign of $\alpha\mapsto c_{1/2}(\alpha)$ would therefore allow one to weaken the definition of an admissible curvilinear polygon without changing the proofs.
We conjecture that $c_{1/2}(\alpha)>0$ for all $\alpha\in(0,\pi)$. In that case, every curvilinear polygon with corner angles in $(0,\pi)$ would be admissible in our sense.
1.2. Circular and annular sectors
We record two model families in which the coefficient of $t^{1/2}$ in (1) admits a closed expression. These formulas give concrete examples of the contribution of curved corners to the heat trace at order $t^{1/2}$.
For a curvilinear polygon $\Omega$, write
(8) $H^{D}_{\Omega}(t)=\cdots+a_{1/2}(\Omega)\,\sqrt{t}+O(t\log t),\qquad a_{1/2}(\Omega)=\frac{1}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds+\sum_{j=1}^{n}\mathcal{C}_{1/2}(\alpha_{j},\kappa_{j,+},\kappa_{j,-}).$
In the examples below, every curved corner is a right angle. We therefore use the special value
(9) $c_{1/2}\left(\frac{\pi}{2}\right)=\frac{1}{16\sqrt{\pi}}.$
Note that if $\alpha_{j}=\pi/2$, then $\cos\alpha_{j}=0$ and $\csc\alpha_{j}=1$, and (2) reduces to
(10) $r_{0}\left(\frac{\pi}{2},\kappa_{+},\kappa_{-}\right)=\frac{1}{2}\sqrt{\kappa_{+}^{2}+\kappa_{-}^{2}}.$
Combining (9) and (10) gives
(11) $\mathcal{C}_{1/2}\left(\frac{\pi}{2},\kappa_{+},\kappa_{-}\right)=\frac{1}{32\sqrt{\pi}}\sqrt{\kappa_{+}^{2}+\kappa_{-}^{2}}.$
Circular sectors. Let $R>0$ and $\alpha\in(0,2\pi)$, and set
$S(R,\alpha)=\{(r,\theta)\in\mathbb{R}^{2}:\ 0<r<R,\ 0<\theta<\alpha\}.$
The boundary of $S(R,\alpha)$ consists of two radial segments and one circular arc. There are three corners: the apex with opening angle $\alpha$, and two right-angle corners at the endpoints of the arc.
The basic geometric quantities are
(12) $|S(R,\alpha)|=\frac{\alpha}{2}R^{2},\qquad|\partial S(R,\alpha)|=(\alpha+2)R,\qquad\int_{\partial S(R,\alpha)}\kappa\,ds=\alpha,\qquad\int_{\partial S(R,\alpha)}\kappa^{2}\,ds=\frac{\alpha}{R}.$
The formula (1) then becomes, after a bit of simplification,
(13) $H^{D}_{S(R,\alpha)}(t)=\frac{\alpha R^{2}}{8\pi t}-\frac{(\alpha+2)R}{8\sqrt{\pi t}}+\frac{1}{12\pi}\Big(\frac{3\pi}{2}+\frac{\pi^{2}+\alpha^{2}}{2\alpha}\Big)+a_{1/2}(S(R,\alpha))\sqrt{t}+O(t\log t).$
To find $a_{1/2}$, observe that at the apex, $(\kappa_{+},\kappa_{-})=(0,0)$, hence $\mathcal{C}_{1/2}=0$. At each endpoint of the circular arc, the corner angle is $\pi/2$ and the limiting curvatures are $(\kappa_{+},\kappa_{-})=(1/R,0)$, so (11) gives a contribution $1/(32\sqrt{\pi}R)$. Summing the two endpoints yields
(14) $\sum_{j}\mathcal{C}_{1/2}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})=\frac{1}{16\sqrt{\pi}}\cdot\frac{1}{R}.$
Substituting (12) and (14) into (8) gives
(15) $a_{1/2}\bigl(S(R,\alpha)\bigr)=\frac{1}{256\sqrt{\pi}}\left(\frac{\alpha}{R}\right)+\frac{1}{16\sqrt{\pi}}\left(\frac{1}{R}\right)=\frac{\alpha+16}{256\sqrt{\pi}}\cdot\frac{1}{R},$
which completes our calculation.
Suppose that two circular sectors are Dirichlet isospectral. Then they must be congruent; their congruence can already be deduced from the coefficients of orders $t^{-1}$, $t^{-1/2}$ and $t^{0}$. We give an alternative proof here to show that the coefficients $a_{-1}$ and $a_{1/2}$ are enough by themselves.
Proposition 3 (Spectral determination within the sector family).
[…]
Proof.
Dirichlet isospectrality implies equality of the heat invariants $|\partial\Omega|$ and $a_{1/2}(\Omega)$. Using (12) and (15), we compute
$|\partial S(R,\alpha)|\cdot 256\sqrt{\pi}\,a_{1/2}\bigl(S(R,\alpha)\bigr)=(\alpha+2)R\cdot\frac{\alpha+16}{R}=(\alpha+2)(\alpha+16).$
Hence $\alpha$ is determined as the unique positive solution of
$x^{2}+18x+32=|\partial\Omega|\cdot 256\sqrt{\pi}\,a_{1/2}(\Omega),$
and then $R=|\partial\Omega|/(\alpha+2)$. Uniqueness of $\alpha$ follows since $x^{2}+18x+32$ is strictly increasing on $(0,\infty)$. ∎
Annular sectors. Let
$A(r,R,\alpha)=\{(\rho,\theta)\in\mathbb{R}^{2}:\ r<\rho<R,\ 0<\theta<\alpha\}.$
Then
(16) $|A(r,R,\alpha)|=\frac{\alpha}{2}(R^{2}-r^{2}),\qquad|\partial A(r,R,\alpha)|=\alpha(R+r)+2(R-r).$
[…]
Thus we have
(17) $H^{D}_{A(r,R,\alpha)}(t)=\frac{\alpha(R^{2}-r^{2})}{8\pi t}-\frac{\alpha(R+r)+2(R-r)}{8\sqrt{\pi t}}+\frac{1}{4}+a_{1/2}(A(r,R,\alpha))\sqrt{t}+O(t\log t).$
Using (16),
(18) $a_{1/2}\bigl(A(r,R,\alpha)\bigr)=\frac{\alpha+16}{256\sqrt{\pi}}\left(\frac{1}{R}+\frac{1}{r}\right).$
This additional coefficient may be used to distinguish annuli which have equal heat coefficients up to that point. For example, set
$\Omega_{1}=A\left(1,3,\frac{1}{4}\right),\qquad\Omega_{2}=A\left(2,4,\frac{1}{6}\right).$
A direct check gives
$|\Omega_{1}|=|\Omega_{2}|=1,\qquad|\partial\Omega_{1}|=|\partial\Omega_{2}|=5,$
so $\Omega_{1}$ and $\Omega_{2}$ have the same heat coefficients through $t^{0}$ within this model family. However,
$a_{1/2}(\Omega_{1})=\frac{16+\frac{1}{4}}{256\sqrt{\pi}}\left(1+\frac{1}{3}\right)=\frac{65}{768\sqrt{\pi}},\qquad a_{1/2}(\Omega_{2})=\frac{16+\frac{1}{6}}{256\sqrt{\pi}}\left(\frac{1}{2}+\frac{1}{4}\right)=\frac{97}{2048\sqrt{\pi}},$
so $a_{1/2}(\Omega_{1})\neq a_{1/2}(\Omega_{2})$. Hence $a_{1/2}$ distinguishes noncongruent annular sectors that are indistinguishable by the coefficients through order $t^{0}$.
This said, even after including $a_{1/2}$, the data $(A=|\Omega|,P=|\partial\Omega|,a_{1/2})$ can arise from two distinct triples $(r,R,\alpha)$, hence from two noncongruent annular sectors. To see this, $s=R+r$ and d=Rrd=R-r. We have $A=\frac{\alpha}{2}ds$ and $P=\alpha s+2d$, so eliminating $\alpha$ gives
(19) $P=\frac{2A}{d}+2d.$
For fixed $(A,P)$ with $P>4\sqrt{A}$, the quadratic $2d^{2}-Pd+2A=0$ has two positive roots, so $d$ is not uniquely determined by $(A,P)$. Moreover, from (18) we deduce
(20) $\bigl(256\sqrt{\pi}\,a_{1/2}(\Omega)\bigr)\,(s^{2}-d^{2})=64s+\frac{8A}{d}.$
For each fixed $d>0$, this is a quadratic equation in $s$ with one positive root and one negative root. Thus each admissible choice of $d$ determines a unique $s>d$, and hence a unique triple $(r,R,\alpha)$.
Therefore, even after incorporating the curved-corner invariant in Theorem 2, the data $(A,P,a_{1/2})$ can correspond to two noncongruent annular sectors. To break this remaining ambiguity one would need additional spectral information, for instance higher heat coefficients or another spectral invariant such as the first Dirichlet eigenvalue.
1.3. Plan of the paper
In section 2 we prove Theorem 1. This is done by examining the geometric microlocal description of the heat kernel given in [10]. We show that there is a well-defined corner contribution $\mathcal{C}_{1/2}\sqrt{t}$, and then prove that any two corners with the same angle $\alpha$ and limit curvatures $\kappa_{\pm}$ yield the same contribution, regardless of the rest of $\Omega$. To take advantage of this uniformity, in section 3 we introduce a simple conformal model for a curvilinear corner and compute the contribution $\mathcal{C}_{1/2}\sqrt{t}$ in that model. This computation proceeds via a parametrix construction as in [10] and allows us to obtain both the factorization (3) and the trace formula (4). Finally, in section 4, we prove (5), i.e. compute $c_{1/2}(\pi/2)$, by computing a sufficient number of terms in the Dirichlet heat trace expansion for a unit half-disk and comparing with (1). Section 4 also contains the same calculation for the Neumann case, identifying $\mathcal{C}^{N}_{1/2}\bigl(\frac{\pi}{2},1,0\bigl)$.
Acknowledgements
D. S. is grateful for support from the AMS-Simons research enhancement grant 501949-9208. S. L. was supported by an Olga Taussky and John Todd Fellowship and by NSF grant DMS-2346799.
We thank Javier Gómez-Serrano for suggesting the problem to us.
2. The corner contribution
Our approach to understanding the heat kernel on $\Omega$ is based on the methods of geometric microlocal analysis [9, 10]. The strategy is to first resolve the singularities of the heat kernel at $t=0$ by constructing a “double heat space,” $\Omega_{h}^{2}$, via a sequence of geometric blow-ups. The heat kernel lifts under these blow-ups to a well-behaved (“polyhomogeneous conormal”) function on the manifold with corners $\Omega_{h}^{2}$ [10]. Roughly speaking, polyhomogeneous conormal functions are those whose asymptotic expansions at the boundary hypersurfaces are compatible and jointly polyhomogeneous at higher-codimension corners; the precise definition may be found in [9].
The boundary of $\Omega_{h}^{2}$ consists of boundary hypersurfaces corresponding to distinct asymptotic regimes. Among these hypersurfaces are:
  • The interior diagonal (td), modeling heat flow far from the boundary.
  • The side faces (sf), modeling heat flow near a smooth edge.
  • The front faces (ff), modeling heat flow at a corner.
The method of proof in [10], which is important for our work here, is to use geometric microlocal analysis to perform a parametrix construction. Specifically, one constructs a parametrix $K(t,z,z^{\prime})$, an approximate solution to the heat equation, by specifying its leading order behavior at each of the three boundary hypersurfaces td, sf, and ff. This should be done in such a way so that when the heat operator is applied to the parametrix $K(t,z,z^{\prime})$, the result $R(t,z,z^{\prime})$ is lower order at td, sf, and ff. In order to arrange this, we choose leading order behavior that solves the relevant model problems for the heat equation at each boundary hypersurface. We also must ensure that the specified behaviors at td, sf, and ff are consistent with each other at the intersections of the boundary hypersurfaces, so that there actually exists a polyhomogeneous conormal function on $\Omega_{h}^{2}$ with all of the correct behavior. Once we have a good enough parametrix, we can recover the true heat kernel via the Volterra series
$H^{D}_{\Omega}=K-K\star R+K\star R\star R-\cdots.$
To analyze this Volterra series, we use a composition formula for polyhomogeneous conormal functions on $\Omega_{h}^{2}$, which in our setting is [10, Theorem 3.16].
To be specific about what we mean when we talk about orders: set $T=\sqrt{t}$. We measure orders at each of td, sf, and ff in powers of $T$; a function which has leading order behavior $T^{-2}$ has order 2-2. “Lower order” means larger powers of $T$. Throughout, “vanishes to infinite order at td (or sf, or ff)” means the expansion has no terms of any finite $T$-order there.
The model cone heat kernel has order 2-2 at td, sf, and ff. We write $\operatorname{ord}_{\mathrm{ff}}(\cdot)$ for this $T$-order. The key result of [10, Theorem 3.16] for our setting is that if $B$ vanishes to infinite order at td, then $\operatorname{ord}_{\mathrm{ff}}(A\star B)=\operatorname{ord}_{\mathrm{ff}}(A)+\operatorname{ord}_{\mathrm{ff}}(B)+4$.
As a first use of these techniques, we show that the corner contribution $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})$ indeed depends only on $\alpha$ and $\kappa_{\pm}$. Suppose that there exist two domains $\Omega$ and $\Omega^{\prime}$, each with one corner, with the same $\alpha$ and the same $\kappa_{\pm}$. By [10], the heat kernels $H^{D}_{\Omega}$ and $H^{D}_{\Omega^{\prime}}$ are polyhomogeneous conormal on $\Omega_{h}^{2}$ and $(\Omega^{\prime})_{h}^{2}$ respectively. Since $\alpha$ is the same, the faces ff of these respective heat spaces are geometrically identical. We now state a lemma which informally says that: under these conditions, $H^{D}_{\Omega}-H^{D}_{\Omega^{\prime}}$ has leading order at worst $t^{0}$ at ff.
Lemma 1.
Let $\Omega$ and $\Omega^{\prime}$ be domains with a single corner having the same opening angle $\alpha$ and the same one‑sided boundary curvatures $\kappa_{\pm}$ at that corner. Then $H^{D}_{\Omega}-H^{D}_{\Omega^{\prime}}$ has order $\leq T^{0}$ (at worst $T^{0}=t^{0}$) at ff. Equivalently, their ff coefficients of orders $T^{-2}=t^{-1}$ and $T^{-1}=t^{-1/2}$ agree.
Proof.
We use $H^{D}_{\Omega^{\prime}}$ to build a parametrix $K$ for $H^{D}_{\Omega}$. We will build this parametrix $K$ for $H^{D}_{\Omega}$ by prescribing its leading behavior at the three faces. Specifically:
  • At td, we specify that $K$ should have the usual interior expansion, which in this setting is simply the Euclidean heat kernel on $\mathbb{R}^{2}$, which has order 2-2 with no lower order terms.
  • At sf, we specify that $K$ should have the same expansion as the Dirichlet heat kernel on a smooth domain with boundary whose boundary coincides with the appropriate smooth part of $\partial\Omega$. Such a heat kernel is known to have a polyhomogeneous conormal expansion by [6], and that expansion is consistent with the Euclidean heat kernel at td. The first term, at order 2-2, is universal. The second term, at order 1-1, follows from the boundary calculus construction in [6]; by a standard scaling argument, this coefficient is a universal multiple of the curvature $\kappa(z^{\prime})$.
  • At ff, we specify that $K$ should have the same $t^{-1}$ (order 2-2) and $t^{-1/2}$ (order 1-1) terms as $H^{D}_{\Omega^{\prime}}$. These are consistent with the interior expansion at td, since they are consistent for $H^{D}_{\Omega^{\prime}}$.
[…]
This consistency means that there exists a parametrix $K(t,z,z^{\prime})$ with the specified expansions at td, sf, and ff. This parametrix solves the model problems to infinite order at td and sf and to two orders at ff. So $R(t,z,z^{\prime})=(\partial_{t}-\Delta_{\Omega})K(t,z,z^{\prime})$ has leading order 2-2, as opposed to 4-4, at ff, and leading order $\infty$ at td and sf. Using the composition formula of [10], we conclude that for each $n\geq 1$, $K\star(R^{\star n})$ has leading order at worst $t^{0}$ at ff.
$\operatorname{ord}_{\mathrm{ff}}(K\star R^{\star n})\geq(-2)+n(-2)+4n=2n-2\geq 0\quad(n\geq 1).$
Thus
$H^{D}_{\Omega}-K=\sum_{n\geq 1}(-1)^{n}K\star R^{\star n}$
has order at worst 0 at ff, and since the same is true for $H^{D}_{\Omega^{\prime}}-K$ by construction of $K$, so $H^{D}_{\Omega}-H^{D}_{\Omega^{\prime}}=O(T^{0})$ at ff. The Lemma is proven. ∎
We can now prove Theorem 1.
Proof of Theorem 1.
As in [10], the heat trace is obtained by restricting $H^{D}_{\Omega}$ to the spatial diagonal and then integrating over $z\in\Omega$. The technical tool used is Melrose’s pushforward theorem. We will not repeat all the details here, but the upshot is that each face td, sf, and ff gives a separate contribution to the expansion (1). A priori, the expansions can interact to give logarithmic terms, but for the same reasons as in [10, p. 49], keeping in mind that we now have two orders at ff rather than one, there are no logarithmic terms until at least $O(T^{2}\log T)=O(t\log t)$ (and potentially much later, if at all). The contribution from td is
$\frac{|\Omega|}{4\pi t}+O(t^{\infty}).$
The contributions from sf are
$\frac{|\partial\Omega|}{8\sqrt{\pi t}}+\frac{1}{12\pi}\int_{\partial\Omega}\kappa\,ds+\frac{\sqrt{t}}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds+O(t).$
And the contributions from ff, since the first is already known and the second depends only on $\alpha$ and $\kappa_{\pm}$, are
$\sum_{j=1}^{n}\frac{\pi^{2}-\alpha_{j}^{2}}{24\pi\alpha_{j}}+\sqrt{t}\sum_{j=1}^{n}\mathcal{C}_{1/2}(\alpha_{j},\kappa_{j,+},\kappa_{j,-})+O(t),$
for some unknown function $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})$. Summing the three faces gives Theorem 1. ∎
Everything in these proofs works in the Neumann case as well.
Note that we will perform a similar parametrix construction, in detail, in the next section, when we build a conformal model to find an expression for $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})$.
3. The conformal model and the construction
The goal of this section is to find an expression for the contribution $\mathcal{C}_{1/2}$. Since it depends only on $\alpha$ and on $\kappa_{\pm}$, we can work with a specific domain with those values. We take advantage of a conformal transformation to build such a domain, and then we construct its heat kernel.
3.1. The conformal model domain
We define a conformal mapping from $\mathbb{C}$ to $\mathbb{C}$. Let $w$ be the complex coordinate on the target space and $z$ be the complex coordinate on the source space. The Euclidean metric on the target $\mathbb{C}$ is $g_{Euc}=dw\,d\bar{w}$. Fix $z_{0}\in\mathbb{C}$ and let $\Phi=\Phi_{z_{0}}:\mathbb{C}\to\mathbb{C}$ be the holomorphic map given by
$w=\Phi(z)=z+z_{0}z^{2}.$
Near the origin, it is a small perturbation of the identity map.
To compute the pullback metric $\Phi^{*}(g_{Euc})$, we use the standard fact that for a holomorphic map $\Phi$, $dw=\Phi^{\prime}(z)dz$ and $d\bar{w}=\overline{\Phi^{\prime}(z)}d\bar{z}$. Therefore,
$\Phi^{*}(g_{Euc})$ $=\Phi^{*}(dw\,d\bar{w})$
$=(\Phi^{\prime}(z)dz)(\overline{\Phi^{\prime}(z)}d\bar{z})$
$=\Phi^{\prime}(z)\overline{\Phi^{\prime}(z)}dz\,d\bar{z}$
$=|\Phi^{\prime}(z)|^{2}dz\,d\bar{z}$
Thus, the pullback metric is:
$\Phi^{*}(g_{Euc})=|1+2z_{0}z|^{2}dz\,d\bar{z}.$
Let $F(z)=|1+2z_{0}z|^{2}$ be the conformal factor derived above. An immediate calculation yields
(21) $F(z)=1+4\Re(z_{0}z)+4|z_{0}|^{2}|z|^{2}.$
If we write $z=x+iy$ and $z_{0}=u_{0}+iv_{0}$, we obtain
(22) $F(x+iy)=1+4(u_{0}x-v_{0}y)+4(u_{0}^{2}+v_{0}^{2})(x^{2}+y^{2}).$
On the other hand, if we write $z=re^{i\theta}$ and $z_{0}=r_{0}e^{i\theta_{0}}$ (this $r_{0}$ matches (2); see (26), below), we obtain
(23) $F(re^{i\theta})=1+4r_{0}r\cos(\theta+\theta_{0})+4r_{0}^{2}r^{2}.$
We see that $F$, however it is written, is smooth, though of course not holomorphic, from $\mathbb{C}\to\mathbb{R}$. It is also smooth down to zero in polar coordinates $(r,\theta)$.
For later use, we also record the gradient of $F$ at the origin:
(24) $\nabla F(0,0)=\langle 4u_{0},-4v_{0}\rangle=4r_{0}\langle\cos\theta_{0},-\sin\theta_{0}\rangle,$
where the latter expression is in polar coordinates but still measuring $x$ and $y$-derivatives,
We now revert to $\mathbb{R}^{2}$. Let $\Omega_{0}$ be the flat sector $\{z\in\mathbb{C}\mid 0<\arg(z)<\alpha\}$. The map $\Phi$ takes $\Omega_{0}$ to a curvilinear domain $\Omega$ with the same corner angle $\alpha$. The signed curvatures of the two boundary edges of $\Omega$ at the corner are given by:
  • $\kappa_{+}$ (on the image of the $\theta=0$ ray): $2\Im(z_{0})$
  • $\kappa_{-}$ (on the image of the $\theta=\alpha$ ray): $2\Im(z_{0}e^{i\alpha})$
Here the orientation is chosen so that the signed curvatures measure curvature in the inward-pointing direction, so that a convex domain will have $\kappa_{+}\geq 0$ and $\kappa_{-}\geq 0$. A domain with symmetry about $\arg(z)=\alpha/2$ will have $\kappa_{+}=\kappa_{-}$.
By choosing the complex parameter $z_{0}$, we can match any desired curvatures $\kappa_{+},\kappa_{-}$, as long as $\alpha\neq\pi$. Solve the following linear system for $z_{0}=u_{0}+iv_{0}$:
$\kappa_{+}$ $=2v_{0}$
$\kappa_{-}$ $=-2\Im((u_{0}+iv_{0})(\cos\alpha+i\sin\alpha))=-2(u_{0}\sin\alpha+v_{0}\cos\alpha)$
This yields $z_{0}=u_{0}+iv_{0}$, where
(25) $\begin{cases}u_{0}=&-\frac{1}{2}\kappa_{+}\cot\alpha-\frac{1}{2}\kappa_{-}\csc\alpha;\\ v_{0}=&\frac{1}{2}\kappa_{+}.\end{cases}$
In terms of the curvatures, we observe that
(26) $r_{0}=\frac{1}{2}|\csc\alpha|\sqrt{\kappa_{+}^{2}+2\kappa_{+}\kappa_{-}\cos\alpha+\kappa_{-}^{2}};\quad\theta_{0}=\textrm{arg}(u_{0}+iv_{0}).$
The pullback by $\Phi$ of the Euclidean metric on $\Omega$ is
$F\cdot g_{Eucl,\Omega_{0}},$
where we now interpret $F$ as a function with domain $\mathbb{R}^{2}$ rather than $\mathbb{C}$. By the formula for the Laplacian in two dimensions under a conformal change of metric $g=Fg_{0}$, the new Laplacian is $\Delta_{g}=F^{-1}\Delta_{g_{0}}$. Therefore, the pullback Laplacian is
$\Phi^{*}(\Delta_{\Omega})=F^{-1}\Delta_{\Omega_{0}}.$
We want to understand the heat kernel for the operator $\Delta_{\Omega}$ on the curved domain $\Omega$. By pulling back, this is equivalent to studying the heat kernel for the operator $F^{-1}\Delta_{\Omega_{0}}$ on $\Omega_{0}$. The corresponding heat equation is $\partial_{t}H=F^{-1}\Delta_{\Omega_{0}}H$. We define our heat operator on the model space $\Omega_{0}$ as:
$\mathcal{L}=\partial_{t}-F^{-1}\Delta_{\Omega_{0}}.$
We now construct the heat kernel for $\mathcal{L}$. We use the techniques of [10], working at td, then at sf and ff. This corresponds to building a parametrix for the heat kernel, first at the interior diagonal, and then at the diagonal both near the boundary and near a corner. Once we have a sufficiently accurate parametrix, we iterate away the error via a Volterra series to produce the true heat kernel.
3.2. The heat kernel at td
It is proved in [10] that the heat kernel for $\mathcal{L}$, being the pullback of the heat kernel for the Laplacian on the curvilinear polygon $\Omega$, is polyhomogeneous conormal on the double space of [10]. We also know that its full expansion at td is the same as that for the heat kernel of the operator $\Phi^{*}(\Delta_{\mathbb{R}^{2}})=F^{-1}\Delta_{\mathbb{R}^{2}}$. This expansion may be written in terms of the (interior) boundary defining function $T=\sqrt{t}$ as
$T^{-2}a_{-2}+T^{-1}a_{-1}+T^{0}a_{0}+\ldots,$
where $a_{-2}$, $a_{-1}$, $a_{0}$, etc. are smooth functions on the face td. We refer to $a_{k}$ as the td-coefficients of the heat kernel; in particular, $a_{0}$ denotes the coefficient of order $T^{0}$ at td, whose restriction along $\mathrm{td}\cap\mathrm{ff}$ will determine $\mathcal{C}_{1/2}(\alpha,\kappa_{+},\kappa_{-})$.
[…]
We now use the work of Grieser [6] to compute $a_{-2}$ and $a_{-1}$. We use his notation throughout this subsection: the coordinates are $(t,x,y)$, where $x\in\mathbb{R}^{2}$ is an interior coordinate on the first factor and $y\in\mathbb{R}^{2}$ on the second. The initial parametrix for the heat kernel is
(27) $K_{1}(t,x,y)=\frac{1}{4\pi t}\exp\left(-\frac{F(y)|x-y|^{2}}{4t}\right).$
Computing the application of the heat operator $(\partial_{t}-F^{-1}(x)\Delta_{\mathbb{R}^{2}})$ to $K_{1}$, we obtain the error term
(28) $R_{1}(t,x,y):=(\partial_{t}-F^{-1}(x)\Delta_{\mathbb{R}^{2}})K_{1}(t,x,y)=\left(1-\frac{F(y)}{F(x)}\right)\frac{\partial}{\partial t}K_{1}(t,x,y).$
Notation (td symbols). In the rescaled variables $(t,X,y)$ with $X=(x-y)/\sqrt{t}$, any heat kernel $A$ will admit an asymptotic of the form
(29) $A(t,x,y)\sim\sum_{q\in d-\frac{1}{2}\mathbb{N}_{0}}t^{-2-q}\,\Phi_{q}(A)(X,y)$
as $t\downarrow 0$ with $X$ fixed. The coefficient $\Phi_{q}(A)$ is the td-symbol of order $q$ of the operator $A$; if $A$ has order $d$, then $\Phi_{d}(A)$ is the principal td-symbol of $A$. Referring to Grieser [6] for the definition of the various operator spaces $\Psi_{H}^{d}$, we have $K_{1}(t,x,y)\in\Psi_{H}^{-1}$. Roughly, this means that $K_{1}(t,x,y)$ has an expansion of the form (29) with d=1d=-1, hence leading term of order $t^{-1}$, and with coefficients decaying to infinite order at $|X|=\infty$.
The prefactor is $O(\sqrt{t})$ near the diagonal. Since $K_{1}\in\Psi_{H}^{-1}$, one has $\partial_{t}K_{1}\in\Psi_{H}^{0}$, which has size $t^{-2}$, so $R_{1}(t,y+\sqrt{t}\,X,y)$ has size $t^{-3/2}$ in td coordinates. Therefore $R_{1}\in\Psi_{H}^{-1/2}$ with principal td-symbol
(30) $\Phi_{-3/2}(R_{1})(X,y)=(\frac{X\cdot\nabla(F)(y)}{F(y)})(\frac{F(y)}{4}X^{2}-1)\frac{1}{4\pi}\exp(-\frac{F(y)}{4}X^{2}).$
By the heat kernel construction in Grieser [6], the leading and subleading terms of the true heat kernel at td will be the same as those of
$K_{1}-K_{1}*R_{1}.$
$K_{1}$ has zero subleading term in the coordinates $(t,X,y)$, so for the subleading term we just need to understand $K_{1}*R_{1}$, in particular just the leading term of $K_{1}*R_{1}$, which is $\Phi_{-3/2}(K_{1}*R_{1})$. This can be computed using [6, Proposition 2.6]. Observe that the integration in $Z$ is with respect to F(y)dZF(y)\,dZ, the metric at $y$. We obtain, including the minus sign,
(31) $\Phi_{-3/2}(-K_{1}*R_{1})(X,y)=-\int_{0}^{1}\int_{\mathbb{R}^{2}}(1-\sigma)^{-1}\sigma^{-3/2}\frac{1}{4\pi}\exp(-\frac{F(y)}{4}\frac{(X-Z)^{2}}{1-\sigma})\\ \frac{Z\cdot\nabla(F)(y)}{\sqrt{\sigma}F(y)}(\frac{F(y)}{4}\frac{Z^{2}}{\sigma}-1)\frac{1}{4\pi}\exp(-\frac{F(y)}{4}\frac{Z^{2}}{\sigma})\,F(y)dZ\,d\sigma.$
This integral can be computed, which we now do. Let
$c=\frac{F(y)}{4};\quad V=\nabla(F)(y);\quad\tilde{Z}=\frac{Z}{\sqrt{\sigma}};\quad\tilde{X}=\frac{X}{\sqrt{\sigma}}.$
We obtain
(32) $\Phi_{-3/2}(-K_{1}*R_{1})=-\frac{1}{16\pi^{2}}\int_{0}^{1}(1-\sigma)^{-1}\sigma^{-1/2}\\ \int_{\mathbb{R}^{2}}\tilde{Z}\cdot V(c\tilde{Z}^{2}-1)\exp(-c\tilde{Z}^{2}-c\sigma\frac{(\tilde{X}-\tilde{Z})^{2}}{1-\sigma})\,d\tilde{Z}\,d\sigma.$
As usual, we can complete the square to do the inner integral. Returning to $X$, we get
$\exp(-cX^{2})\int_{\mathbb{R}^{2}}\tilde{Z}\cdot V(c\tilde{Z}^{2}-1)\exp(-\frac{c}{1-\sigma}(\tilde{Z}-\sqrt{\sigma}X)^{2})\,d\tilde{Z}$.
Let $\hat{Z}=\tilde{Z}-\sqrt{\sigma}X$:
$\exp(-cX^{2})\int_{\mathbb{R}^{2}}(\hat{Z}+\sqrt{\sigma}X)\cdot V(c(\hat{Z}+\sqrt{\sigma}X)^{2}-1)\exp(-\frac{c}{1-\sigma}\hat{Z}^{2})\,d\hat{Z}$.
Any of the terms in the prefactor
$((\hat{Z}+\sqrt{\sigma}X)\cdot V)(c(\hat{Z}+\sqrt{\sigma}X)^{2}-1)$
which are odd in one of the coordinates of $\hat{Z}$ yield zero when integrated. Up to such terms, that pre-factor is
$(\sqrt{\sigma}X\cdot V)(c\hat{Z}^{2}+c\sigma X^{2}-1)+2c\sqrt{\sigma}(\hat{z}_{1}^{2}x_{1}v_{1}+\hat{z}_{2}^{2}x_{2}v_{2}).$
Using the usual formulas for Gaussian integrals, we get that the inner integral in (32) is
$\exp(-cX^{2})[(\sqrt{\sigma}X\cdot V)(c\sigma X^{2}-1)\frac{\pi(1-\sigma)}{c}+(\sqrt{\sigma}X\cdot V)\frac{2\pi(1-\sigma)^{2}}{c}]$
$c^{-1}\pi\sigma^{1/2}(1-\sigma)\exp(-cX^{2})(X\cdot V)(c\sigma X^{2}+1-2\sigma),$
and therefore
(33) $\Phi_{-3/2}(-K_{1}*R_{1})=-\frac{1}{16c\pi}\exp(-cX^{2})(X\cdot V)\int_{0}^{1}(c\sigma X^{2}+1-2\sigma)\,d\sigma,$
Performing the simple integration, we obtain
(34) $\Phi_{-3/2}(-K_{1}*R_{1})=-\frac{1}{32\pi}\exp(-cX^{2})(X\cdot V)X^{2}.$
Inserting the variables, we have
(35) $\Phi_{-3/2}(-K_{1}*R_{1})=-\frac{1}{32\pi}X^{2}(X\cdot\nabla F(y))\exp(-\frac{F(y)}{4}X^{2}).$
We can now write down the two leading orders of the heat kernel at td: they are
(36) $T^{-2}\frac{1}{4\pi}\exp(-\frac{F(y)}{4}X^{2})-T^{-1}\frac{1}{32\pi}X^{2}(X\cdot\nabla F(y))\exp(-\frac{F(y)}{4}X^{2}).$
As one may check, this expression is in the kernel of $T^{2}\mathcal{L}$ up to $O(T^{0})$.
Remark.
Two observations:
  • If $F$ is constant, the subleading term is zero. In fact all subleading terms are zero – this is the Euclidean setting, where the initial parametrix is equal to the true heat kernel.
  • In all cases, $\Phi_{-3/2}(K_{1}*R_{1})(0,y)=0$. This means that the restriction of $K_{1}*R_{1}$ to the spatial diagonal has no term of order $T^{-1}$ at td. This is unsurprising, as we do not expect any term of the form $T^{-1}=t^{-1/2}$ in the interior heat trace expansion.
3.3. A parametrix for the heat kernel
We take our heat operator $\mathcal{L}$ and apply it to a parametrix
$H^{(1)}:=G\cdot H_{cone},$
where $H_{cone}$ is the Dirichlet heat kernel for an exact sector of angle $\alpha$, and where $G$ is a scalar function chosen as a corrector to allow $H^{(1)}$ to satisfy the heat equation to two orders at td.
Notation. In this subsection we set $T=\sqrt{t}$ and use the rescaled difference vector
$X:=\frac{z-z^{\prime}}{T}\in\mathbb{R}^{2},\qquad|X|=\frac{|z-z^{\prime}|}{T}.$
We write $X\cdot\nabla F(z^{\prime})$ for the Euclidean dot product and $|X|^{2}=X\cdot X$.
Specifically, define
(37) $G:=\Bigl(1-\frac{1}{8}\,T\,|X|^{2}\,(X\cdot\nabla F(z^{\prime}))\Bigr)\exp\Bigl[-\frac{F(z^{\prime})-1}{4}\,|X|^{2}\Bigr].$
Then $G$ times the Euclidean heat kernel (the td model for $H_{\operatorname{cone}}$) yields (36). We do not claim that $G$ is polyhomogeneous on the double space; if $F(z^{\prime})<1$ then the factor $\exp\bigl(\frac{1-F(z^{\prime})}{4}|X|^{2}\bigr)$ grows as $|X|\to\infty$ (note that $|X|=|z-z^{\prime}|/T$). However, $G$ is polyhomogeneous away from $\mathrm{td}$, and $G\,H_{cone}$ is polyhomogeneous on the double space. Moreover, $G\,H_{cone}$ satisfies Dirichlet boundary conditions [^2] The Neumann problem is harder precisely because such a parametrix would not necessarily inherit Neumann boundary conditions. This can likely be worked around but we do not do so here., inheriting this from $H_{cone}$.
Now examine $G$ at ff, the corner front face. This face is the most singular stratum of the heat space, where $t\to 0$ and both points $x,y=0$. The function $G$ is equal to 1 at ff because $T=0$ at ff, and $F=1$ at ff. As a consequence, $H^{(1)}$ is in the kernel of the heat operator $\mathcal{L}$ to leading order at ff; we say that it “solves the model problem” to one order at ff. It also solves the model problem to two orders at td. And since $T=0$ at sf and $\exp[-\frac{1}{4}(F(y)-1)X^{2}]H_{cone}$ agrees to first order at sf with the initial parametrix $\exp[-\frac{1}{4}(F(y)-1)X^{2}]H_{\mathbb{R}^{2}_{+}}$ used by Grieser at sf [6], $H^{(1)}$ also solves the model problem to one order at sf.
Now define:
$R^{(1)}:=\mathcal{L}(H^{(1)}).$
If $H^{(1)}$ were not correctly chosen, we would expect $R^{(1)}$ to be two orders worse than $H^{(1)}$ at each of td, ff, and sf, which means order 4-4 at each. However, since $H^{(1)}$ solves the model problem to two orders at td and one order at ff, and also one order at sf, $R^{(1)}$ has leading orders given by 2-2 at td, 3-3 at ff, and 3-3 at sf.
In order to use the composition theorem of [10, Theorem 3.16], which requires one of the two operators to vanish to infinite order at td, we work with a slightly improved parametrix $H^{(2)}$ given by solving away the full expansion at td. This equals $H^{(1)}$ plus a term which is order zero at td, ff, and sf. Let $D:=H^{(2)}-H^{(1)}$. Then $D$ has order $\geq 0$ at ff. Hence if we use the notation
$R^{(2)}:=\mathcal{L}(H^{(2)}),$
then $R^{(2)}-R^{(1)}=\mathcal{L}(D)$ has order $\geq-2$ at ff and sf. Moreover, $R^{(2)}$ vanishes to infinite order at td and it has leading order 3-3 at both ff and sf.
By Duhamel’s principle and our usual Neumann series sum, the true heat kernel for our conformal metric is
$H=H^{(2)}-H^{(2)}*R^{(2)}+\sum_{k=2}^{\infty}(-1)^{k}H^{(2)}*(R^{(2)})^{k}.$
We are interested in finding its term of order 1-1 at ff (the sub-leading term). By the composition formula in [10], using in particular that
$\text{ord}_{\text{ff}}(A*B)=\text{ord}_{\text{ff}}(A)+\text{ord}_{\text{ff}}(B)+4,$
the terms with $k\geq 2$ have no order 1-1 term at ff, so the order 1-1 term of our true heat kernel at ff is the same as that of
$H^{(2)}-H^{(2)}*R^{(2)}.$
This in turn [^3] The justification for using the simpler parametrix $H^{(1)}$ to compute the subleading term at the ff face is done by an analysis of operator orders within the composition calculus of [10]. Let $D=H^{(2)}-H^{(1)}$ be the correction term that improves the parametrix at the td face. By construction, the leading order of $D$ at the ff face is at least $T^{0}$. The difference between the first two terms of the respective Neumann series is $\Delta H=D-H^{(1)}*\mathcal{L}(D)-D*R^{(1)}-D*\mathcal{L}(D)$. A direct application of the composition formula shows that each term in this expression has a leading order of $T^{0}$ or higher at the ff face. For instance, the most singular term, $H^{(1)}*\mathcal{L}(D)$, has an order of at least 2+(2)+4=0-2+(-2)+4=0. Since the entire difference $\Delta H$ is no more singular than $T^{0}$, the expressions $H^{(1)}-H^{(1)}*R^{(1)}$ and $H^{(2)}-H^{(2)}*R^{(2)}$ must agree for all terms of order $T^{-1}$ and below. We are therefore justified in using the simpler expression to compute the desired subleading coefficient. is the same as that of
$H^{(1)}-H^{(1)}*R^{(1)}=GH_{cone}-GH_{cone}*(\mathcal{L}(GH_{cone})).$
Here $H^{(1)}$ has order 2-2 at ff, so we need its leading and subleading terms. $R^{(1)}$ has leading ff order 3-3. Thus the leading ff term of the convolution $H^{(1)}*R^{(1)}$ comes from $T^{-2}\ast T^{-3}$, which lands at 1-1 at ff, by the ff composition rule.
A critical fact here is that by the composition formula in [10], the leading term of $H^{(1)}*R^{(1)}$ at ff depends only on the leading terms of $H^{(1)}$ and $R^{(1)}$ at ff, not on their leading terms at other boundary faces (in particular at sf). This is helpful for the following reasons.
  • First, the integral in the convolution $H^{(1)}*R^{(1)}$ is done with respect to the metric $Fg_{Eucl}$, but since we are computing at the ff face, which corresponds to the vertex where $r=0$, we can evaluate $F(z)$ at that point. As established, $F(0)=1$. So, for the purpose of calculating the leading term of the convolution, the integration measure simplifies from the Riemannian volume form to the standard Euclidean one.
  • Second, when computing the second convolution factor $\mathcal{L}(GH_{cone})$, we can ignore any term below its leading order 3-3. Similarly, for the first convolution factor, $GH_{\text{cone}}$, we only need its leading order term, which is of order $T^{-2}$. All smoother terms can be ignored as they will not affect the leading term of the result.
This leads to the following simplification. Consider $GH_{cone}*(\mathcal{L}(GH_{cone}))$. We only need to worry about the leading order of each term, so $G$ may be replaced with 1 and we need only analyze
$H_{cone}*(\mathcal{L}(GH_{cone})).$
Let us compute $\mathcal{L}(GH_{cone})$. It is
(38) $\mathcal{L}(GH_{cone})=\mathcal{L}(G)H_{cone}+G\mathcal{L}(H_{cone})-2F^{-1}\nabla G\cdot\nabla(H_{cone}).$
Using the fact that $\mathcal{L}_{cone}H_{cone}=0,$ this becomes
(39) $\mathcal{L}(GH_{cone})=\mathcal{L}(G)H_{cone}-G(F^{-1}-1)\Delta(H_{cone})-2F^{-1}\nabla G\cdot\nabla(H_{cone}).$
We may replace the $\mathcal{L}$ with $\mathcal{L}_{cone}$ in the first term, as $(\mathcal{L}-\mathcal{L}_{cone})(G)$ has order zero at ff. We may also drop the $G$ in the second term and the $F^{-1}$ in the third term, as they are both 1 to leading order at ff. Thus up to terms which are irrelevant for our calculation,
$\mathcal{L}(GH_{cone})\sim\mathcal{L}_{cone}(G)H_{cone}-(F^{-1}-1)\Delta(H_{cone})-2\nabla G\cdot\nabla(H_{cone}).$
In conclusion, to two orders at ff, we have
(40) $H\sim GH_{cone}-H_{cone}*(\mathcal{L}_{cone}(G)H_{cone}-(F^{-1}-1)\Delta(H_{cone})-2\nabla G\cdot\nabla H_{cone}),$
where the convolution – and all derivatives here – are taken with respect to the Euclidean metric on the cone. The geometric correction factor $G$ must be retained in the stand-alone factor $G\,H_{\mathrm{cone}}$ (before taking the trace): its $T^{+1}$ contribution times the $T^{-2}$ of $H_{\mathrm{cone}}$ produces one of the two $T^{-1}$ pieces in the kernel at ff we are calculating.
3.4. Computations with $G$
We now turn to the explicit calculation of the terms in Equation (40). We are interested in the first two terms in their expansion at ff.
Throughout, we replace Grieser’s $x$ and $y$ by $z$ and $z^{\prime}$. Let $z$ have polar coordinates $(r,\theta)$ and let $z^{\prime}$ have polar coordinates $(r^{\prime},\theta^{\prime})$, and then let
$R=r/T,R^{\prime}=r^{\prime}/T.$
With this notation, natural local coordinates near the interior of the front face ff are $(T,R,\theta,R^{\prime},\theta^{\prime})$, where $T$ is a boundary defining function. In these coordinates, we compute the first two terms in the expansion (40), of orders $T^{-2}$ and $T^{-1}$. To do this we compute two terms in the expansion of $GH_{cone}$. We also compute the leading $T^{-2}$ term in the expansion of $H_{cone}$ and the leading $T^{-3}$ term in the expansion of the second convolution factor, as by the composition formula in [10] their convolution will give a term of order $T^{-2-3+4}=T^{-1}$ at ff. To this end, we denote
(41) $a(R,\theta,R^{\prime},\theta^{\prime})=T^{2}H_{cone}|_{\mathrm{ff}};$
(42) $b(R,\theta,R^{\prime},\theta^{\prime})=r_{0}^{-1}(T^{3}(\mathcal{L}_{cone}(G)H_{cone}-(F^{-1}-1)\Delta(H_{cone})-2\nabla G\cdot\nabla H_{cone}))|_{\mathrm{ff}}.$
Our goal is to find expressions for $a$ and $b$.
Let $\varphi_{j}(\theta)$ be the $L^{2}$-normalized eigenfunction
$\varphi_{j}(\theta)=\sqrt{\frac{2}{\alpha}}\sin(\frac{\pi j\theta}{\alpha}).$
First, we use Cheeger’s formula for the Dirichlet heat kernel on the exact sector of angle $\alpha$ [2]. In polar coordinates $z=(r,\theta)$ and $z^{\prime}=(r^{\prime},\theta^{\prime})$, one has
$H_{\mathrm{cone}}(t;r,\theta,r^{\prime},\theta^{\prime})=\frac{1}{2t}\exp\Big[-\frac{1}{4t}\big(r^{2}+(r^{\prime})^{2}\big)\Big]\sum_{j=1}^{\infty}I_{\mu_{j}}\Big(\frac{rr^{\prime}}{2t}\Big)\,\varphi_{j}(\theta)\,\varphi_{j}(\theta^{\prime}).$
Now set $T=\sqrt{t}$, $R=r/T$, and $R^{\prime}=r^{\prime}/T$. Since $r=TR$, $r^{\prime}=TR^{\prime}$, and $t=T^{2}$, this becomes $H_{\mathrm{cone}}=T^{-2}a(R,\theta,R^{\prime},\theta^{\prime})$ by a simple scaling argument, with $a$ given by
(43) $a(R,\theta,R^{\prime},\theta^{\prime})=\frac{1}{2}\,\exp\Big[-\frac{R^{2}+(R^{\prime})^{2}}{4}\Big]\sum_{j=1}^{\infty}I_{\mu_{j}}\Big(\frac{RR^{\prime}}{2}\Big)\,\varphi_{j}(\theta)\,\varphi_{j}(\theta^{\prime}).$
To be absolutely clear, we write
$H_{\mathrm{cone}}(t;r,\theta,r^{\prime},\theta^{\prime})=T^{-2}\,a\left(\frac{r}{\sqrt{t}},\theta,\frac{r^{\prime}}{\sqrt{t}},\theta^{\prime}\right),\qquad T=\sqrt{t},$
For all other terms we must analyze $G$. Recall that $G$ is a function of $T=\sqrt{t}$, the source point $z^{\prime}$, and the scaled, relative position vector $X=(z-z^{\prime})/T$. We have the relations
$z$ $=(r\cos\theta,r\sin\theta)=(RT\cos\theta,RT\sin\theta)$
$z^{\prime}$ $=(r^{\prime}\cos\theta^{\prime},r^{\prime}\sin\theta^{\prime})=(R^{\prime}T\cos\theta^{\prime},R^{\prime}T\sin\theta^{\prime}).$
We can directly compute the components of the scaled vector $X$:
(44) $X=(R\cos\theta-R^{\prime}\cos\theta^{\prime},R\sin\theta-R^{\prime}\sin\theta^{\prime}).$
Define the distance function $D=D(R,\theta,R^{\prime},\theta^{\prime})$ by
(45) $D^{2}:=|X|^{2}=(R\cos\theta-R^{\prime}\cos\theta^{\prime})^{2}+(R\sin\theta-R^{\prime}\sin\theta^{\prime})^{2}=R^{2}+(R^{\prime})^{2}-2RR^{\prime}\cos(\theta-\theta^{\prime}).$
For future use we also define
(46) $S=S(R,\theta,R^{\prime},\theta^{\prime})=\left(R\cos(\theta+\theta_{0})+R^{\prime}\cos(\theta^{\prime}+\theta_{0})\right).$
With these relations, we can write the full expression for $G$ in the front-face coordinates:
(47) $G(T,R,\theta,R^{\prime},\theta^{\prime})=\Bigg[1-\frac{1}{8}TD^{2}\langle R\cos\theta-R^{\prime}\cos\theta^{\prime},R\sin\theta-R^{\prime}\sin\theta^{\prime}\rangle\cdot\nabla F(R^{\prime}T,\theta^{\prime})\Bigg]\\ \exp[-\frac{1}{4}(F(R^{\prime}T,\theta^{\prime})-1)D^{2}].$
Now we compute the expansion of $G$ at ff in powers of $T$. Since $F$ is smooth, we have by Taylor series that
$G\approx 1-\frac{1}{8}TD^{2}\langle R\cos\theta-R^{\prime}\cos\theta^{\prime},R\sin\theta-R^{\prime}\sin\theta^{\prime}\rangle\cdot\nabla F(0,0)-\frac{1}{4}TD^{2}\langle R^{\prime}\cos\theta^{\prime},R^{\prime}\sin\theta^{\prime}\rangle\cdot\nabla F(0,0)+O(T^{2}).$
By combining the vector coefficients and plugging in $\nabla F(0,0)=4r_{0}\langle\cos\theta_{0},-\sin\theta_{0}\rangle$, this expression simplifies:
(48) $G$ $=1-TD^{2}\left[\frac{1}{8}\langle R\cos\theta+R^{\prime}\cos\theta^{\prime},R\sin\theta+R^{\prime}\sin\theta^{\prime}\rangle\right]\cdot\left(4r_{0}\langle\cos\theta_{0},-\sin\theta_{0}\rangle\right)+O(T^{2})$
(49) $=1-\frac{r_{0}}{2}TD^{2}\langle R\cos\theta+R^{\prime}\cos\theta^{\prime},R\sin\theta+R^{\prime}\sin\theta^{\prime}\rangle\cdot\langle\cos\theta_{0},-\sin\theta_{0}\rangle+O(T^{2})$
(50) $=1-\frac{r_{0}}{2}TD^{2}\left(R(\cos\theta\cos\theta_{0}-\sin\theta\sin\theta_{0})+R^{\prime}(\cos\theta^{\prime}\cos\theta_{0}-\sin\theta^{\prime}\sin\theta_{0})\right)+O(T^{2})$
(51) $=1-\frac{r_{0}}{2}TD^{2}S+O(T^{2}).$
From this we immediately conclude that, at ff,
(52) $GH_{cone}=T^{-2}a(R,\theta,R^{\prime},\theta^{\prime})-\frac{r_{0}}{2}T^{-1}D^{2}Sa(R,\theta,R^{\prime},\theta^{\prime})+O(1).$
One nice feature of this expression is that, since $D^{2}$ is zero on the spatial diagonal, the sub-leading term of $GH_{cone}$ gives no contribution to the heat trace.
The last step is to find an expression for $b(R,\theta,R^{\prime},\theta^{\prime})$ by simplifying (42). As a first step, we examine the middle term in (42). We have
$F^{-1}(r,\theta)=1-4r_{0}r\cos(\theta+\theta_{0})+O(r^{2}),$
so in our front face coordinates,
$(F^{-1}-1)=-4r_{0}TR\cos(\theta+\theta_{0})+O(T^{2}).$
We use the notation
$\Delta_{R,\theta}=\partial_{R}^{2}+\frac{1}{R}\partial_{R}+\frac{1}{R^{2}}\partial_{\theta}^{2},\qquad\nabla_{R,\theta}=(\partial_{R},\frac{1}{R}\partial_{\theta}).$
With this notation, we have $\nabla_{r,\theta}=T^{-1}\nabla_{R,\theta}$ and $\Delta_{r,\theta}=T^{-2}\Delta_{R,\theta}$, and so
$\Delta H_{cone}=T^{-2}\Delta_{R,\theta}H_{cone},$
which by conic scaling is
$T^{-4}\Delta_{R,\theta}a(R,\theta,R^{\prime},\theta^{\prime}).$
Therefore
$(F^{-1}-1)\Delta H_{cone}=4r_{0}T^{-3}R\cos(\theta+\theta_{0})\Delta_{R,\theta}a(R,\theta,R^{\prime},\theta^{\prime}),$
and $T^{3}r_{0}^{-1}$ times this is the contribution that is made to $b(R,\theta,R^{\prime},\theta^{\prime})$:
(53) $T^{3}r_{0}^{-1}(F^{-1}-1)\Delta H_{cone}=4R\cos(\theta+\theta_{0})\Delta_{R,\theta}a.$
The other terms in (42) involve derivatives of $G$, so we must compute these using (48). At ff,
$\nabla G=-\frac{r_{0}}{2}\nabla_{R,\theta}(D^{2}S)+O(T);\quad\Delta G=-\frac{r_{0}}{2}T^{-1}\Delta_{R,\theta}(D^{2}S)+O(1).$
On the other hand, $\partial_{t}$ lifts to
$\frac{1}{2}T^{-2}(T\partial_{T}-R\partial_{R}-R^{\prime}\partial_{R^{\prime}})$
and so
(54) $\partial_{t}G=-\frac{a_{0}}{4}T^{-1}D^{2}S+\frac{a_{0}}{4}T^{-1}(R\partial_{R}+R^{\prime}\partial_{R}^{\prime})(D^{2}S)+O(1).$
All in all,
(55) $\mathcal{L}G=-\frac{a_{0}}{4}T^{-1}(1-R\partial_{R}-R^{\prime}\partial_{R^{\prime}}-2\Delta_{R,\theta})(D^{2}S)+O(1).$
Lemma 2.
With $\Delta_{R,\theta}$ and $\nabla_{R,\theta}$ the polar Laplacian/gradient in $(R,\theta)$,
$\Delta_{R,\theta}(D^{2})=4,\qquad\Delta_{R,\theta}S=0,\qquad\Delta_{R,\theta}(D^{2}S)=8R\cos(\theta+\theta_{0}),\qquad(R\partial_{R}+R^{\prime}\partial_{R^{\prime}})(D^{2}S)=3D^{2}S.$
Proof.
This is a direct computation, using for instance that $\Delta_{R,\theta}R^{2}=4$ and $S$ is harmonic in $(R,\theta)$. ∎
Putting all of this together, we obtain
(56) $b(R,\theta,R^{\prime},\theta^{\prime})=\left(\frac{1}{2}D^{2}S+4R\cos(\theta+\theta_{0})\right)a(R,\theta,R^{\prime},\theta^{\prime})\\ +4R\cos(\theta+\theta_{0})\Delta_{R,\theta}a(R,\theta,R^{\prime},\theta^{\prime})\\ +\nabla_{R,\theta}(D^{2}S)\cdot\nabla_{R,\theta}a(R,\theta,R^{\prime},\theta^{\prime}).$
For later use, we rewrite the gradient term by using the identity
(57) $\nabla u\cdot\nabla v=\frac{1}{2}\Delta(uv)-\frac{1}{2}v\Delta u-\frac{1}{2}u\Delta v.$
This yields, after some computations,
(58) $b(R,\theta,R^{\prime},\theta^{\prime})$ $=\Big(\tfrac{1}{2}D^{2}S\Big)\,a(R,\theta,R^{\prime},\theta^{\prime})$
+ $\quad+4R\cos(\theta+\theta_{0})\,\Delta_{R,\theta}a(R,\theta,R^{\prime},\theta^{\prime})$
$\quad+\,\tfrac{1}{2}\,\Delta_{R,\theta}\big(D^{2}S\cdot a(R,\theta,R^{\prime},\theta^{\prime})\big)-\tfrac{1}{2}\,(D^{2}S)\,\Delta_{R,\theta}a(R,\theta,R^{\prime},\theta^{\prime}).$
Either (56) or (58) may be used.
3.4.1. The little trace and the convolution
To find the heat trace, we first restrict our kernel to the diagonal to find the so-called “little trace”. By (52), the restriction to the diagonal of the $T^{-1}$ coefficient of $GH_{cone}$ is zero, and therefore the first term in (40) does not contribute to the $T^{-1}$ term of interest. The only contribution comes from the second term in (40), and that contribution is the convolution of the leading orders of the two factors. This convolution is
$\int_{0}^{t}\int_{0}^{\infty}\int_{0}^{\alpha}(t-s)^{-1}a(\frac{r}{\sqrt{t-s}},\theta,\frac{r^{\prime}}{\sqrt{t-s}},\theta^{\prime})s^{-3/2}b(\frac{r^{\prime}}{\sqrt{s}},\theta^{\prime},\frac{r^{\prime\prime}}{\sqrt{s}},\theta^{\prime\prime})\,r^{\prime}\,d\theta^{\prime}\,dr^{\prime}\,ds.$
Switching to front face variables yields that the term of interest, $H_{\text{sub}}$, is, where $\cong$ denotes having the same heat trace coefficient at order $T^{-1}$,
(59) $H_{\text{sub}}\cong-T^{-1}\int_{0}^{1}\int_{0}^{\infty}\int_{0}^{\alpha}(1-\sigma)^{-1}\sigma^{-3/2}\\ \times a\left(\frac{R}{\sqrt{1-\sigma}},\theta,\frac{R^{\prime}}{\sqrt{1-\sigma}},\theta^{\prime}\right)b\left(\frac{R^{\prime}}{\sqrt{\sigma}},\theta^{\prime},\frac{R^{\prime\prime}}{\sqrt{\sigma}},\theta^{\prime\prime}\right)\,R^{\prime}\,d\theta^{\prime}\,dR^{\prime}\,d\sigma.$
Setting $(R^{\prime\prime},\theta^{\prime\prime})=(R,\theta)$ gives the little trace. If you also make the change of variables $\tilde{R}^{\prime}=R^{\prime}/\sqrt{\sigma}$, then $R^{\prime}dR^{\prime}=\sigma\tilde{R}^{\prime}d\tilde{R}^{\prime}$. Then
(60) $(\textrm{tr }H_{\text{sub}})(R,\theta)=-T^{-1}\int_{0}^{1}\int_{0}^{\infty}\int_{0}^{\alpha}(1-\sigma)^{-1}\sigma^{-1/2}\\ \times a\left(\frac{R}{\sqrt{1-\sigma}},\theta,\frac{\sqrt{\sigma}\tilde{R}^{\prime}}{\sqrt{1-\sigma}},\theta^{\prime}\right)b\left(\tilde{R}^{\prime},\theta^{\prime},\frac{R}{\sqrt{\sigma}},\theta\right)\,\tilde{R}^{\prime}\,d\theta^{\prime}\,d\tilde{R}^{\prime}\,d\sigma.$
To find the contribution to the heat trace, we first restrict this kernel to the diagonal by setting $(R^{\prime},\theta^{\prime})=(R,\theta)$, then we integrate with respect to $r\,dr\,d\theta=T^{2}R\,dR\,d\theta$. By Melrose’s pushforward theorem (see [9]; see also [5, Theorem 3.10] for a precise statement), the trace contribution is given by a renormalized integral. Writing down this renormalized integral yields (4).
The last thing to do is to address the dependence of this expression on $\kappa_{\pm}$. By (58), the kernel $b$ does appear to depend on $\theta_{0}$ through factors such as $\cos(\theta+\theta_{0})$. However, integrating in the angular variables removes this dependence. Let $I(\theta_{0})$ denote the right-hand side of (4). From (58), the $\theta_{0}$ dependence enters only through shifted angles of the form $\theta+\theta_{0}$ and $\phi+\theta_{0}$, while the remaining angular dependence, coming from the $D^{2}$ term, occurs through differences such as $\theta-\phi$. Differentiating under the integral sign therefore yields
$\frac{d}{d\theta_{0}}I(\theta_{0})=\iint_{(0,\alpha)\times(0,\alpha)}(\partial_{\theta}+\partial_{\phi})(\cdots)\,d\theta\,d\phi,$
where $(\cdots)$ denotes the $\theta,\phi$ dependent integrand in (4). Integrating by parts in $\theta$ and $\phi$ reduces this to boundary terms at $\theta,\phi\in\{0,\alpha\}$, and these vanish for Dirichlet boundary conditions since the sector kernel $a$ in (43) contains the factors $\varphi_{j}(\theta)$ and $\varphi_{j}(\phi)$ with $\varphi_{j}(0)=\varphi_{j}(\alpha)=0$. Hence $\frac{d}{d\theta_{0}}I(\theta_{0})=0$, so $c_{1/2}(\alpha)=I(\theta_{0})$ is independent of $\theta_{0}$. This justifies writing $c_{1/2}(\alpha)$, rather than $c_{1/2}(\alpha,\kappa_{-},\kappa_{+})$, on the left-hand side of (4).
4. The case $\alpha=\pi/2$
4.1. The Dirichlet case
Our goal here is to compute through the $t^{1/2}$ term in the short-time expansion of the heat trace of the Laplacian on a domain $\Omega$ which is a half-disk of radius 1, with Dirichlet boundary conditions. This provides a concrete test case in which the corner term can be read off from an explicit spectrum and compared directly with (1) and (3). This heat trace is denoted
(61) $\mathrm{Tr}\ e^{-t\Delta_{\Omega}}.$
The spectrum of the Dirichlet Laplacian on $\Omega$ consists of $\{j_{m,k}^{2}\}_{m,k\in\mathbb{N}}$, each with multiplicity one. This may be compared to the spectrum of the Dirichlet Laplacian on a disk $D$, which consists of all of these eigenvalues but with multiplicity two, as well as the set $\{j_{0,k}^{2}\}_{k\in\mathbb{N}}$, each with multiplicity one. As a result, if we let
(62) $H(t):=\sum_{k=1}^{\infty}e^{-tj_{0,k}^{2}},$
then
(63) $\mathrm{Tr}\ e^{-t\Delta_{\Omega}}=\frac{1}{2}\mathrm{Tr}\ e^{-t\Delta_{D}}-\frac{1}{2}H(t).$
We will take advantage of this.
The heat trace $\mathrm{Tr}\ e^{-t\Delta_{D}}$ is well understood and has a full polyhomogeneous asymptotic expansion in half-integer powers as $t\to 0$. The beginning of this expansion is, from [4],
(64) $\mathrm{Tr}\ e^{-t\Delta_{D}}=\frac{1}{4}t^{-1}-\frac{\sqrt{\pi}}{4}t^{-1/2}+\frac{1}{6}+\frac{\sqrt{\pi}}{128}t^{1/2}+O(t).$
For $H(t)$, we use the asymptotics of Bessel function zeroes due to McMahon [3, (10.21.19)]:
(65) $j_{0,k}=(k-\frac{1}{4})\pi+\frac{1}{8(k-\frac{1}{4})\pi}+O(k^{-3}),$
which imply
(66) $j_{0,k}^{2}=(k-\frac{1}{4})^{2}\pi^{2}+\frac{1}{4}+O(k^{-2}).$
Based on this we define the function
(67) $\tilde{H}(t)=\sum_{k=1}^{\infty}\exp[-t((k-\frac{1}{4})^{2}\pi^{2}+\frac{1}{4})]$
and estimate its difference with $H(t)$.
Lemma 3.
With all notation as above, as $t\to 0$,
(68) $H(t)-\tilde{H}(t)=O(t).$
This is helpful because we also have the following lemma.
Lemma 4.
As $t\to 0$,
(69) $\tilde{H}(t)=\frac{1}{2\sqrt{\pi}}t^{-1/2}+c-\frac{1}{8\sqrt{\pi}}t^{1/2}+O(t).$
Here $c$ is a constant which is irrelevant to our purposes.
From these two Lemmas and (63) we immediately deduce this proposition.
Proposition 4.
As $t\to 0$,
(70) $\mathrm{Tr}\ e^{-t\Delta_{\Omega}}=\frac{1}{8}t^{-1}-(\frac{\sqrt{\pi}}{8}+\frac{1}{4\sqrt{\pi}})t^{-1/2}-\frac{1}{2}c+(\frac{\sqrt{\pi}}{256}+\frac{1}{16\sqrt{\pi}})t^{1/2}+O(t).$
Remark.
We could reverse engineer $c$ from the known expansion from [10] but there is no need.
We are now in a position to state what the contribution of each corner of the half-disk to the $t^{1/2}$ term is. The $t^{1/2}$ term above is the sum of a curvature integral over the boundary and a pair of identical corner contributions. From [4] we see that the $\frac{\sqrt{\pi}}{256}$ is precisely the curvature integral. Dividing the remainder in half shows
Theorem 3.
The contribution to the $t^{1/2}$ coefficient in the heat trace of a half-disk of radius 1 from each of its two corners is
(71) $\frac{1}{32\sqrt{\pi}}t^{1/2}.$
Combined with (3), and the fact that $r_{0}(\frac{\pi}{2},1,0)=\frac{1}{2}$, this identifies $c_{1/2}(\pi/2)$ and confirms that right-angled curved corners contribute nontrivially at order $t^{1/2}$. It remains only to prove the Lemmas.
4.2. Proofs of the Lemmas
First we prove Lemma 3.
Proof.
Rearrangement of the various expressions gives
(72) $H(t)-\tilde{H}(t)=e^{-t/4}\sum_{k=1}^{\infty}e^{-t(k-\frac{1}{4})^{2}\pi^{2}}(1-e^{-t\cdot O(k^{-2})}).$
Here $O(k^{-2})$ is shorthand for $j_{0,k}^{2}-(k-\frac{1}{4})^{2}\pi^{2}-\frac{1}{4}$, which is bounded in absolute value by $Ck^{-2}$ for some universal constant $C$. Since $k\geq 1$ and $t\leq 1$ is all that is relevant, we have that for $t\leq 1$, $|1-e^{-t\cdot O(k^{-2})}|\leq\tilde{C}tk^{-2}$ for a universal constant $\tilde{C}$. This implies
(73) $|H(t)-\tilde{H}(t)|\leq\tilde{C}t\sum_{k=1}^{\infty}e^{-t(k-\frac{1}{4})^{2}\pi^{2}}k^{-2}.$
Estimating each exponential from above by 1 yields that the right-hand side is $O(t)$ as desired. ∎
Now we prove Lemma 4.
Proof.
First write
(74) $\tilde{H}(t)=e^{-t/4}\hat{H}(t),$
where
(75) $\hat{H}(t)=\sum_{k=1}^{\infty}\exp[-t(k-\frac{1}{4})^{2}\pi^{2}].$
We know from [10] that $\mathrm{Tr}\ e^{-t\Delta_{\Omega}}$ has a full polyhomogeneous asymptotic expansion, and therefore by (63), so does $H(t)$. By Lemma 3, $\tilde{H}(t)$ has at least a polyhomogeneous expansion up to $O(t)$, and therefore so does $\hat{H}(t)$. The upshot of this is that one may define a zeta function
(76) $\hat{\zeta}(s)=\frac{1}{\Gamma(s)}\int_{0}^{\infty}\hat{H}(t)t^{s-1}\,dt,$
and that this zeta function, initially defined for $\Re(s)>>0$, has a meromorphic continuation at least to the half-plane $\Re(s)>-1$. Moreover, by the usual argument (agreement for $\Re(s)>>0$),
(77) $\hat{\zeta}(s)=\sum_{k=1}^{\infty}((k-\frac{1}{4})\pi)^{-2s}.$
But now we recognize the zeta function $\hat{\zeta}(s)$ as a form of a Hurwitz zeta function. Specifically,
(78) $\hat{\zeta}(s)=\pi^{-2s}\zeta(2s,\frac{3}{4}),$
where $\zeta(s,a)$ is the Hurwitz zeta function. The function $\zeta(s,a)$ has a single, simple pole at $s=1$, with residue 1 [3, 25.11(i)]. So $\hat{\zeta}(s)$ has a single, simple pole at $s=1/2$, and its residue there is $\frac{1}{2\pi}$. Thus the function
(79) $\int_{0}^{\infty}\hat{H}(t)t^{s-1}\,dt=\Gamma(s)\hat{\zeta}(s)$
has poles at $s=1/2$ and also at each non-positive integer; moreover its residue at $s=1/2$ is $\Gamma(1/2)/(2\pi)=\frac{1}{2\sqrt{\pi}}$.
We can now reverse-engineer the expansion of $\hat{H}(t)$ as $t\to 0$, which is morally taking an inverse Mellin transform. We already know such a partial expansion exists up to order $O(t)$, and now we know that its coefficients are at orders $t^{-1/2}$, $t^{0}$, and higher integer powers of $t$ which are absorbed in $O(t)$. The coefficient of $t^{-1/2}$ is precisely the residue of $\Gamma(s)\hat{\zeta}(s)$ at $1/2$, which is $\frac{1}{2\sqrt{\pi}}$. So, for some constant $c$ which is not relevant for our purposes,
(80) $\hat{H}(t)=\frac{1}{2\sqrt{\pi}}t^{-1/2}+c+O(t).$
Multiplying by $e^{-t/4}\sim 1-\frac{t}{4}+O(t^{2})$,
(81) $\tilde{H}(t)=\frac{1}{2\sqrt{\pi}}t^{-1/2}+c-\frac{1}{8\sqrt{\pi}}t^{1/2}+O(t),$
completing the proof. ∎
4.3. A Neumann right-angle corner contribution
This section computes the order-$t^{1/2}$ contribution coming from the two right-angle corners of the unit half-disk with Neumann boundary conditions. The computation uses the explicit Neumann spectrum on the disk and half-disk, together with the smooth-boundary coefficient for Neumann boundary conditions. The final outcome is an explicit value for the corner remainder in this model geometry, which we record as $\mathcal{C}^{N}_{1/2}\bigl(\frac{\pi}{2},1,0\bigr)$.
Let $\Omega\subset\mathbb{R}^{2}$ be the unit half-disk
$\Omega=\{(r,\theta):0<r<1,\ 0<\theta<\pi\},$
endowed with Neumann boundary conditions along both the diameter and the semicircle. Let $D$ denote the unit disk, also with Neumann boundary conditions, and write $\Delta^{N}_{\Omega}$ and $\Delta^{N}_{D}$ for the corresponding Laplacians. For the unit disk with Neumann boundary conditions, the smooth-boundary heat trace expansion gives (see Proposition 8 of [4])
(82) $\mathrm{Tr}\ e^{-t\Delta^{N}_{D}}=\frac{1}{4}t^{-1}+\frac{\sqrt{\pi}}{4}t^{-1/2}+\frac{1}{6}+\frac{5\sqrt{\pi}}{128}t^{1/2}+O(t).$
Separation of variables shows that the Neumann spectrum on the disk is
$\mathrm{spec}(\Delta_{D}^{N})=\{0\}\ \cup\ \{(j^{\prime}_{m,k})^{2}:\ m\in\mathbb{N}_{0},\ k\in\mathbb{N}\},$
where $j^{\prime}_{m,k}$ is the $k$th positive zero of $J_{m}^{\prime}$, with multiplicity $1$ for $m=0$ and multiplicity $2$ for $m\geq 1$ (the $\cos(m\theta)$ and $\sin(m\theta)$ modes). On the half-disk, the Neumann condition on $\theta=0,\pi$ forces $\partial_{\theta}u=0$, hence only the $\cos(m\theta)$ modes occur, so each $(j^{\prime}_{m,k})^{2}$ has multiplicity $1$. Thus
(83) $\mathrm{Tr}\ e^{-t\Delta^{N}_{\Omega}}=\frac{1}{2}\,\mathrm{Tr}\ e^{-t\Delta^{N}_{D}}+\frac{1}{2}\,H^{N}(t),$
where the correction term is precisely the $m=0$ contribution
$H^{N}(t):=\sum_{\lambda\in\{0\}\cup\{(j^{\prime}_{0,k})^{2}\}_{k\geq 1}}e^{-t\lambda}=1+\sum_{k=1}^{\infty}e^{-t(j^{\prime}_{0,k})^{2}}.$
Using $J_{0}^{\prime}(x)=-J_{1}(x)$, this may be rewritten as
(84) $H^{N}(t)=1+\sum_{k=1}^{\infty}e^{-tj_{1,k}^{2}},$
where $\{j_{1,k}\}_{k\geq 1}$ are the positive zeros of $J_{1}$.
McMahon’s asymptotics for Bessel zeros (cf. [3, (10.21.19)]) give
(85) $j_{1,k}=\Bigl(k+\frac{1}{4}\Bigr)\pi-\frac{3}{8\bigl(k+\frac{1}{4}\bigr)\pi}+O(k^{-3}),\qquad k\to\infty,$
hence
(86) $j_{1,k}^{2}=\Bigl(k+\frac{1}{4}\Bigr)^{2}\pi^{2}-\frac{3}{4}+O(k^{-2}).$
Define the model series
(87) $\widetilde{H}^{N}(t):=1+\sum_{k=1}^{\infty}\exp\Bigl[-t\Bigl(\Bigl(k+\frac{1}{4}\Bigr)^{2}\pi^{2}-\frac{3}{4}\Bigr)\Bigr].$
Lemma 5.
As $t\to 0$ one has $H^{N}(t)-\widetilde{H}^{N}(t)=O(t)$.
Proof.
This is identical to the proof of Lemma 3. ∎
Lemma 6.
As $t\to 0$,
(88) $\widetilde{H}^{N}(t)=\frac{1}{2\sqrt{\pi}}\,t^{-1/2}+c_{N}+\frac{3}{8\sqrt{\pi}}\,t^{1/2}+O(t),$
for some constant $c_{N}\in\mathbb{R}$.
Proof.
Write $\widetilde{H}^{N}(t)=e^{3t/4}\,\widehat{H}^{N}(t)$ with
$\widehat{H}^{N}(t):=e^{-3t/4}\widetilde{H}^{N}(t)=e^{-3t/4}+\sum_{k=1}^{\infty}\exp\Bigl[-t\Bigl(k+\frac{1}{4}\Bigr)^{2}\pi^{2}\Bigr].$
As in the proof of Lemma 4, the existence of a polyhomogeneous expansion for $\widehat{H}^{N}(t)$ through $O(t)$ allows one to define the Mellin transform
$\widehat{\zeta}_{N}(s)=\frac{1}{\Gamma(s)}\int_{0}^{\infty}\widehat{H}^{N}(t)\,t^{s-1}\,dt=\sum_{k=0}^{\infty}\bigl(\pi(k+\tfrac{5}{4})\bigr)^{-2s}=\pi^{-2s}\zeta(2s,\tfrac{5}{4}),$
initially for $\Re(s)>>0$ and then by meromorphic continuation.
Since $\zeta(2s,\tfrac{5}{4})$ has a simple pole at $2s=1$ with residue $1$, the function $\widehat{\zeta}_{N}$ has a simple pole at $s=\tfrac{1}{2}$ with residue $1/(2\pi)$. It follows that $\int_{0}^{\infty}\widehat{H}^{N}(t)t^{s-1}\,dt=\Gamma(s)\widehat{\zeta}_{N}(s)$ has residue $\Gamma(1/2)/(2\pi)=1/(2\sqrt{\pi})$ at $s=1/2$, and hence
$\widehat{H}^{N}(t)=\frac{1}{2\sqrt{\pi}}\,t^{-1/2}+c_{N}+O(t).$
Multiplying by $e^{3t/4}=1+\frac{3}{4}t+O(t^{2})$ yields (88). ∎
Combining Lemmas 5 and 6 shows that $H^{N}(t)$ itself satisfies (88) (with a possibly different constant term), up to $O(t)$. Substituting (82) and (88) into (83) gives the half-disk expansion
(89) $\mathrm{Tr}\ e^{-t\Delta^{N}_{\Omega}}=\frac{1}{8}\,t^{-1}+\Bigl(\frac{\sqrt{\pi}}{8}+\frac{1}{4\sqrt{\pi}}\Bigr)t^{-1/2}+\text{\emph{(const)}}+\Bigl(\frac{5\sqrt{\pi}}{256}+\frac{3}{16\sqrt{\pi}}\Bigr)t^{1/2}+O(t).$
The coefficient of $t^{1/2}$ in (89) may be split into a smooth-boundary term and a corner remainder by subtracting the smooth-boundary contribution for Neumann boundary conditions. For the unit disk, $\int_{\partial D}\kappa^{2}\,ds=2\pi$, and the $t^{1/2}$ coefficient $\frac{5\sqrt{\pi}}{128}$ in (82) corresponds to the smooth-boundary constant $\frac{5}{256\sqrt{\pi}}$. For the half-disk, $\partial\Omega$ consists of a semicircle of curvature $1$ and length $\pi$ and a diameter of curvature 0, so
$\frac{5}{256\sqrt{\pi}}\int_{\partial\Omega}\kappa^{2}\,ds=\frac{5}{256\sqrt{\pi}}\cdot\pi=\frac{5\sqrt{\pi}}{256}.$
Subtracting this from the $t^{1/2}$ coefficient in (89) leaves a remainder $\frac{3}{16\sqrt{\pi}}$. The two endpoints of the semicircle are congruent corners, so they contribute equally. This motivates the definition
(90) $\mathcal{C}^{N}_{1/2}\Bigl(\frac{\pi}{2},1,0\Bigr):=\frac{1}{2}\cdot\frac{3}{16\sqrt{\pi}}=\frac{3}{32\sqrt{\pi}}.$
Here $(\kappa_{+},\kappa_{-})=(1,0)$ records that one incident boundary arc has curvature $1$ (the semicircle) and the other has curvature 0 (the diameter).
Remark.
Equation (90) records the corner value $\mathcal{C}^{N}_{1/2}\bigl(\frac{\pi}{2},1,0\bigr)=\frac{3}{32\sqrt{\pi}}$. If a Neumann factorization of the form
$\mathcal{C}^{N}_{1/2}(\alpha,\kappa_{+},\kappa_{-})=c^{N}_{1/2}(\alpha)\,r_{0}(\alpha,\kappa_{+},\kappa_{-})$
holds for $\alpha=\pi/2$, then since $r_{0}(\pi/2,1,0)=\frac{1}{2}$ by (10), one obtains
$c^{N}_{1/2}\Bigl(\frac{\pi}{2}\Bigr)=\frac{\mathcal{C}^{N}_{1/2}\bigl(\frac{\pi}{2},1,0\bigr)}{r_{0}(\pi/2,1,0)}=\frac{3}{16\sqrt{\pi}}.$